# Query Rewriting via LLMs

### v3

@@ 7.4 Commercial DBMS
Interestingly, although we did not observe any regressions with PostgreSQL, a few did surface in the commercial systems. Nevertheless, our regression identification mechanisms effectively caught these brittle rewrites. As a case in point, a promising rewrite, as estimated by the optimizer, for Q23, actually takes 37 minutes to complete as compared to 18 minutes for the original query – this doubling slowdown was successfully flagged by the SEER heuristic, and the rewrite was abandoned.
We also conducted a preliminary study of LITHE on an internal real-world benchmark with more realistic data and query characteristics than TPC-DS. Even on this benchmark, LITHE produced 8 CPRs (out of 45 queries tested) with CSGM of 2 and TSGM of 3.3.
@@
7.5 Dependence on LLM (Training/Model)
7.5.1 Masked Database
An interesting question to ask now is whether the performance benefits seen thus far could be an artifact of GPT-4o having already been trained well on the TPC-DS benchmark, which is prominent in the public domain. To investigate this issue, we created a masked version of the TPC-DS database schema, whereby the table and column names convey no semantic information about their contents. We then constructed rewrites for the CPR queries (after syntactic changes to reflect the new masked schema) on this version. The results are shown in Table 8, and we observe that the performance profiles only marginally decrease for both LITHE and SOTA, testifying to their robust applicability.
Table 8: Rewrite Performance on Masked Database.
Approach # CPR CSGM TSGM
TPC-DS Masked TPC-DS Masked TPC-DS Masked
LITHE 26 24 11.5 10.5 13.2 11.8
SOTA 13 12 6.1 5.3 4.9 4.2
7.5.2 LITHE on LLaMA
@@
For this environment, Table 9 shows LITHE’s performance on the TPC-DS benchmark with and without MCTS. Although certainly lower than the corresponding numbers with GPT-4o (Section 7.1), it is encouraging to see that, in absolute terms, significant performance benefits can be obtained for most queries, especially with MCTS support. So, the message is that smaller models can also be fruitfully used in real-world environments.
@@ 8.1 Rewrite Space Coverage by LLMs
Given the decades-long research on database query optimization, we expected the potential for performance improvement via rewriting to be limited. What came as a surprise was the substantial scope for improvement still available, as showcased by the large CSGM and TSGM values, even on commercial platforms. These results suggest that LLMs explore optimization spaces that are well outside the purview of contemporary database engines. Further, this enhanced space could be augmented, in a two-stage process, with the recent proposals for LLM-based “plan hints” that steer the optimizer in fruitful directions within a plan space [8].
@@ 8.3 Agentic LLMs for Query Rewriting
An effective way to extend LITHE is to use an agentic LLM that actively interacts with the database environment. This would allow a “LITHE agent” to leverage a memory store to log one or more of: (1) query execution times; (2) memory usages of slow queries; (3) prior interactions with users; and (4) useful/new rewrite rules learnt over time. Additionally, the agent could be given database engine access so that it can directly request metadata such as query execution plans, database statistics, live query analyzers, etc. By following a structured workflow – planning, acting, observing, and refining – the agent would then iteratively enhance query performance. This adaptive workflow supports continuous learning, so that query rewriting can become more efficient over time.
@@ 9 Related Work
LLMs for Database Modules. LLM technologies have been advocated for a variety of database modules. For instance, they have been extensively used for Text-to-SQL transformations [26, 44, 3, 4, 37, 49]. The main focus of these techniques is to correctly ascertain the information necessary to formulate the SQL query [39, 32, 37]. On the other hand, the goal of S2S rewriting is on improving the performance of an existing SQL query. Thus, unlike Text-to-SQL transformations where the input text is inherently ambiguous, SQL queries are precisely defined, and therefore equivalence to a precise ground-truth has to be provably maintained.
In recent times, LLMs have also been considered for plan-hinting [8], join-order optimization [38], index selection [49], data pipelines [22], data management [25], etc. An LLM-enabled multi-modal query optimizer [41], where data spans text, image, audio and video domains, has also been proposed recently. These approaches can be used in conjunction with the LITHE system since they address orthogonal segments of the query processing pipeline.
@@ Appendix A Additional Experiments
A.1 Rewrite Quality (TPC-DS) Contd.
Figure 10 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA for the queries run on a warm-cache environment.
Figure 10: Execution Time Speedups in a warm-cache scenario.
As expected, the speedups often significantly increase when compared to the cold-cache scenario. For instance, the speedup for Q45 increases from 700X to over $10^{4}$. Overall, the TSGM of LITHE increases to 17.1 (from 13.2), while that of SOTA increases to 5.7 (from 4.9).
A.2 Rewrite Quality (Other Benchmarks)
@@ Appendix C Query Equivalence Testing via Sampling
Appendix D Detailed MCTS Algorithm
Algorithm 2 Token-augmented Rewrite
1:$root$,          # Start State;
2:$k$,              # Maximum number of child expansion
3:$\theta$,              # Probability threshold for node expansion
4:$iter_{max}$,     # Maximum number of iterations
5:$Potential$, $visits$, $V$leftarrow empty Map
6:for $i\leftarrow 1,2,\dots,iter_{max}$ do
7:    $visits[root]$leftarrow $visits[root]$ + 1
8:    $n_{cur}$leftarrow $root$
9:    # Stage 1: Selection
10:    while $len(n_{cur}.children)>0$ do
11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n_{cur},a)$
12:        $visits[n_{cur}]$leftarrow $visits[n_{cur}]+1$
13:    # Stage 2: Expansion
14:    $expand\leftarrow$ True
15:    while $expand$ do
16:        $tokens_{next}$, $P_{next}$leftarrow Model($n_{cur},k$)
17:        if $tokens_{next}[0]=$‘;’ then
18:           $expand\leftarrow$ False
19:        else if $P_{next}\leq\theta$ then
20:           for $token\in tokens_{next}$ do
21:               $state\leftarrow n_{cur}.state\cdot token$
22:               $n_{new}\leftarrow$ new $Node$ with State $state$
23:               Append $n_{new}$ to $n_{cur}.children$
24:           $expand\leftarrow$ False
25:        else
26:           $state\leftarrow n_{cur}.state\cdot tokens_{next}[0]$ First Token
27:           $n_{new}\leftarrow$ new $Node$ with State $state$
28:           Append $n_{new}$ to $n_{cur}.children$
29:           $n_{cur}\leftarrow n_{new}$
30:    # Stage 3: Simulation
31:    $query\leftarrow$ GreedyExpand($n_{cur}$)
32:    # Compute the potential of query (returns 0 for invalid query)
33:    $v\leftarrow$ ComputePotential($query$)
34:    $Potential[query]\leftarrow v$
35:    # Stage 4: Backpropagation
36:    $n\leftarrow n_{cur}$
37:    while $n\neq$ Null do
38:        if $v>V[n]$ then
39:           $V[n]\leftarrow v$
40:        $n\leftarrow\text{Parent}(n)$
41:# Return valid rewrite with maximum $Potential>1$
42:if $\exists q\in Potential\mid Potential[q]>1$ then
43:    return $q$ having maximum value of $Potential[q]$
44:else
45:    return the original query
Appendix E Examples used in Prompts for Rules 1–6

### v4

@@ 7.4 Commercial DBMS
Interestingly, although we did not observe any regressions with PostgreSQL, a few did surface in the commercial systems. Nevertheless, our regression identification mechanisms effectively caught these brittle rewrites. As a case in point, a promising rewrite estimated by one of the commercial optimizers for Q23, actually takes 37 minutes to complete as compared to 18 minutes for the original query – this doubling slowdown was successfully flagged by the SEER heuristic (Section 6.3), and the rewrite was abandoned.
@@
7.5 Dependence on LLM
7.5.1 Impact of Training
An interesting question to ask now is whether the performance benefits seen thus far could be an artifact of GPT-4o having already been trained well on the TPC-DS benchmark, which is prominent in the public domain.
To investigate this issue, we ran LITHE on two datasets that were confirmed to be unknown to the GPT-4o version used in our evaluation. The first is the recently released Football benchmark [24] for Text-to-SQL evaluation. From the query suite, which largely consists of simple SPJ and Union queries, we selectively picked the few complex queries. Further, we created our own complex queries to include in the test suite. On this enriched workload, LITHE on PostgreSQL produced 11 CPRs with CSGM of 12 and TSGM of 1.5, while SOTA gave 6 CPRs with CSGM of 3 and TSGM of 1.1.
The second dataset is a proprietary customer benchmark used extensively by an industrial development team. For this environment, LITHE on a commercial engine produced 8 CPRs with a CSGM of 2 and a TSGM of 3.3. Note that these improvements are achieved despite this engine having been fine-tuned on this workload over an extended period. Moreover, the numbers are consistent with those reported for TPC-DS in Section 7.4.
7.5.2 Impact of Model
@@
For this environment, Table 8 shows LITHE’s performance on the TPC-DS queries with and without MCTS. Although certainly lower than the corresponding numbers with GPT-4o (Section 7.1), it is encouraging to see that, in absolute terms, significant performance benefits can be obtained for most queries, especially with MCTS support. So, the message is that smaller models can also be fruitfully used in real-world environments.
@@
Finally, we also evaluated LITHE on Gemini 2.5 Flash. Here, LITHE produced 17 CPR with CSGM of 7.2 and TSGM of 5.8. Although not as strong as GPT-4o, these results highlight the potency of LITHE’s recommendations across a range of LLM platforms.
@@ 8.1 Rewrite Space Coverage by LLMs
Given the decades-long research on database query optimization, we expected the potential for performance improvement via rewriting to be limited. What came as a surprise was the substantial scope for improvement still available, as showcased by the large CSGM and TSGM values, even on commercial platforms. These results suggest that LLMs explore optimization spaces that are well outside the purview of contemporary database engines. Specifically, rewrites in the query space appear to guide the optimizer to explore fresh regions of the plan space not evaluated with the original query formulation. Further, this enhanced space could be augmented, in a two-stage process, with the recent proposals for LLM-based “plan hints” that steer the optimizer in fruitful directions within a plan space [11]. We plan to explore these space relationships and options in our future work.
@@ 8.3 Agentic LLMs for Query Rewriting
An effective way to extend LITHE is to use an agentic LLM that actively interacts with the database environment. This would allow a “LITHE agent” to leverage database-related “tools” such as the query optimizer’s cardinality estimation and cost estimation modules, as well as rewrite-related tools such as rule generation, semantic validation, and regression checking. A memory store could be attached to the agent to log prior interactions with users and rewrite rules learnt over time, thus supporting continuous learning. The agent can decide the order in which these tools are invoked and make appropriate choices during the rewrite process. Finally, combining our MCTS formulation with tool-augmented agentic reasoning would also be a promising avenue for future work.
@@ 8.4 Scope of Semantic Equivalence Tools
8.5 Enhancing NL2SQL approaches with Rewrites
LITHE could potentially be used to enhance NL2SQL approaches to produce performant SQL queries. One way would be to incorporate some rewrite rules directly into NL2SQL prompts. Based on our empirical assessment (Figure 11(b), we would recommend the simplest Prompt 1 from the Basic Prompts in conjunction with a subset of Rules R1–R4. To avoid multiple prompts, one could use the LLM itself to identify the appropriate rule based on the NL query. While these prompts might not always provide the best rewrites, our experiments show that they are sufficient for many scenarios while incurring only small overheads as compared to using the entire prompt suite and MCTS search.
An alternate non-invasive approach would be to simply use LITHE as a post-processor after the initial translation, but again restricted to using only a subset of the prompts/rules to keep the overheads within permissible limits.
@@ 9 Related Work
Query rewriting in SQL engines. Query rewrite prior to plan enumeration has long been a feature of contemporary (non-LLM) database optimizers [9, 25]. However, the difference in our context is that the LLM gives us the flexibility to choose from a wide variety of approaches, not restricted to the hardwired options within individual systems.
LLMs for Database Modules. LLM technologies have been advocated for a variety of database modules. For instance, they have been extensively used for Text-to-SQL transformations [32, 53, 4, 5, 46, 58]. Very recently, statistical metadata was leveraged to also improve the Text-to-SQL generation process [43]. The main focus of these techniques is to correctly ascertain the information necessary to formulate the SQL query [48, 40, 46].
On the other hand, the goal of SQL-to-SQL rewriting is on improving the performance of an existing SQL query. Thus, unlike Text-to-SQL transformations where the input text is inherently ambiguous, SQL queries are precisely defined, and therefore equivalence to a precise ground-truth has to be provably maintained.
In recent times, LLMs have also been considered for a variety of database modules, including plan-hinting [11], join-order optimization [47], index selection [58], data pipelines [28], data management [31], and multi-modal data optimization [50]. These approaches can be used in conjunction with LITHE since they address orthogonal segments of the query processing pipeline.
@@ Appendix A Additional Experiments
A.1 Rewrite Quality (Other Benchmarks)
@@
A.2 Rewrite Quality (Rule-based vs LLM-based approaches)
Our SOTA baseline comprises two theoretical rule-based approaches—Learned Rewrite and LLM-R2—and two LLM-based approaches—GenRewrite and Baseline Prompt. The detailed breakdown of CPR achieved by these methods is presented in Table 10.
Table 10: # CPR of Different SOTA Approaches in Comparison with LITHE
Tools/Techniques Approach Individual # CPR >=>=1.1x
LITHE LLM Based 26 30
Genrewrite LLM Based 11 12
Baseline LLM Based 10 10
Learned Rewrite Rule Based 1 2
LLMR2 Rule Based 0 0
While the theoretical rule-based approaches provide formal correctness guarantees, our TPC-DS experiments revealed that the LLM-based approaches consistently outperformed them. Among the 13 CPRs (i.e., query rewrites with cost speedup $\geq$ 1.5X), only one was produced by a theoretical rule-based approach. Moreover, even when we decreased this threshold to as low as $\geq$ 1.1X, only one more rewrite became a CPR!
LLM-based approaches, on the other hand, seem more suited for query rewrites, resulting in 12 (GenRewrite) and 30 (LITHE) rewrites with a cost speedup $\geq$ 1.1X. This is likely a consequence of the LLM having a “global” vision of the entire query, giving it more leeway in considering rewrites in the query-space. Whereas, theoretical rule-based approaches operate mainly at the operator level within the plan tree, thus being confined to only local optimizations.
A.3 Rewrite quality (Coverage of Basic Prompts)
Figure 11(a) shows the CPR distribution across the Basic Prompts on the Microbenchmark. Among the 6 rewritten queries, 3 are unique to a single prompt, while the other 3 are produced by multiple prompts. At first glance, one might think that Prompt 3 could be removed from our prompt suite since it does not produce any unique rewrite. However, this is an artifact of the micro-benchmark – on the corresponding picture for the full TPC-DS benchmark (Figure 11(b)), Prompt 3 does contribute unique rewrites.
(a) Micro-benchmark
(b) TPC-DS
Figure 11: CPR distribution of Basic Prompts
@@ Appendix B Overheads Reduction for LITHE
As an alternative to the LLM based classifier, we also experimented with a DSPy MIPRO [38] based classifier to identify which, if any, of Rules R1–R6 is appropriate for a given query. However, our initial results indicated that it was weaker than our LLM-based classifier. While the average time taken by this approach reduced to just 10 seconds per query, the CPRs are reduced to 14, and the CSGM and TSGM fall down to 3.3 and 4.5 respectively.
But we hasten to add that this outcome is likely an artifact of our small rule set – looking to the future, where there may be a larger corpus, a data-driven selection method like MIPRO may come to the fore.
@@ Appendix C Query Equivalence Testing via Sampling
Appendix D Examples used in Prompts for Rules 1–6